# Lung-DDPM: Semantic Layout-guided Diffusion Models for Thoracic CT Image Synthesis

### v1

@@ 2.1 Workflow
In this section, we introduce the proposed diffusion model-based thoracic image synthesis method in detail. The workflow of the proposed method is depicted in Figure 2 and consists of four stages:
@@
    Data collection stage: The workflow begins with collecting thoracic CT scans from clinical practices to build a database for case management. At this stage, no cases are annotated.
@@ 2.2 Training and sampling processes
where $m$ is the semantic layout, $z\sim\mathcal{N}(0,\mathbf{I})$, and $\sigma_{t}=\sqrt{\beta_{t}}$. However, the traditional sampling process has a significant drawback: its overdependence on detailed semantic layouts. As demonstrated in Figure 1, without guidance from semantic layouts for extra-pulmonary areas, this process may produce abnormal results with incorrect anatomical structures. In clinical practice, manually annotating every anatomical structure in a thoracic CT scan is impractical. Moreover, semi-automatic and automatic segmentation tools [18, 19] are not sufficiently reliable to provide consistent semantic layouts due to the diverse types and parameter settings of CT scanners [20].
@@
Through the proposed AAS process, we can generate high-quality thoracic CT images guided by radiologist-provided semantic layouts while maintaining anatomically reasonable extra-pulmonary areas, which significantly impact overall image quality. The training and AAS pseudo-codes are depicted in Algorithms 1 and 2, respectively.
@@ 2.3 Network designs
3D U-Net: Inspired by the original 2D U-Net design for diffusion models [17], 3D-DDPM [21] and Med-DDPM [12] introduced the 3D U-Net architecture for brain magnetic resonance imaging (MRI) synthesis. To adapt this network for our thoracic CT image synthesis task, we modified it to accept a noise matrix and the corresponding lung semantic layout, both with dimensions of $128^{3}$. The network input consists of a noised CT image and its corresponding semantic layouts. After passing through a convolution layer, the encoder path (the left part of the 3D U-Net) progressively reduces spatial dimensions while increasing the number of feature channels, enabling the network to learn hierarchical representations of thoracic CT images. The bottleneck (the middle part of the 3D U-Net) processes low-dimensional features to capture global context and relationships within the CT image representations. Finally, the decoder (the right part of the 3D U-Net) reconstructs the spatial resolution of the features, gradually transforming the compact representation back into a full-sized output. Simultaneously, it combines both low-level and high-level features from the encoder via skip connections to generate detailed CT images. The decoder output is the predicted noise $\epsilon^{\prime}$ for the input image.
@@ 2.4 Learning object
During the training stage, we use a pixel-wise least absolute deviations (L1) loss to measure the differences between the added noise $\epsilon$ and the predicted noise $\epsilon^{\prime}$:
@@ 3.1 Description of cohorts
We utilize an internal (Quebec Heart and Lung Institute Research Center (IUCPQ)) dataset and an external dataset (LIDC-IDRI [22]) in this study. The population characteristics of the IUCPQ dataset are summarized in Table 1. Additionally, we provide the patient list from the LIDC-IDRI dataset used in this study on the project’s GitHub page.
Table 1: Study population characteristics of discovery, validation, and test cohorts divided from the IUCPQ dataset. Continuous data were reported as mean ± standard deviation (SD), and categorical data as counts and percentages.
Cohort / Patient group Discovery (Upstream) N=800 Validation N=800 Test N=249
Age, mean (SD) 65.42 (8.26) 65.73 (7.89) 65.04 (8.23)
Sex, N (%)
Male 340 (42.50) 378 (47.25) 105 (42.17)
Female 460 (57.50) 422 (52.75) 144 (57.83)
Race, N (%)
White 731 (91.37) 729 (91.12) 228 (91.57)
Non-white 3 (0.38) 4 (0.50) 1 (0.40)
Unknown 66 (8.25) 67 (8.38) 20 (8.03)
Smoke status, N (%)
Non-smoker 63 (7.87) 57 (7.13) 13 (5.22)
Current smoker 153 (19.13) 167 (20.88) 55 (22.09)
Former smoker 578 (72.25) 571 (71.38) 177 (70.08)
Passive smoker 6 (0.75) 5 (0.63) 4 (1.61)
Nodule count, N (%)
1 712 (89.00) 723 (90.38) 227 (91.16)
$\geq$ 2 88 (11.00) 77 (9.62) 22 (8.84)
Nodule size, mean, mm (SD) 41.59 (30.32) 41.01 (33.90) 41.10 (32.00)
malignant cases N=729 N=744 N=232
Stage, N (%)
0 7 (0.96) 2 (0.27) 2 (0.86)
1 491 (67.54) 511 (68.68) 150 (64.66)
2 123 (16.87) 138 (18.55) 46 (19.83)
3 65 (8.92) 73 (9.81) 20 (8.62)
4 5 (0.69) 5 (0.67) 1 (0.43)
Unknown 38 (5.21) 15 (2.02) 13 (5.60)
Histology, N (%)
Adenocarcinoma 544 (74.62) 535 (71.91) 170 (73.28)
Squamous cell carcinoma 105 (14.40) 122 (16.40) 37 (15.95)
Neuroendocrine tumor 43 (5.90) 47 (6.32) 16 (6.90)
Adenosquamous carcinoma 17 (2.33) 16 (2.15) 5 (2.16)
Sarcomatoid carcinoma 6 (0.82) 12 (1.61) 2 (0.86)
Non-small cell carcinoma 6 (0.82) 6 (0.81) 1 (0.43)
Large cell carcinoma 3 (0.41) 2 (0.27) 0 (0.00)
Mesothelioma 2 (0.27) 0 (0.00) 0 (0.00)
Thymic carcinoma 1 (0.14) 0 (0.00) 0 (0.00)
Other 2 (0.27) 4 (0.54) 1 (0.43)
@@
Table 2: Organization of our datasets (R. indicates real samples, and S. indicates synthetic samples.)
Cohort Patient number Stage
[…]
@@
LIDC-IDRI: To further validate the developed Lung-DDPM under public and transparent experimental conditions, we included the external LIDC-IDRI [22] dataset, developed through collaboration between the Lung Image Database Consortium (LIDC) and the Image Database Resource Initiative (IDRI). This dataset consists of diagnostic and lung cancer screening thoracic CT scans with marked-up annotated lesions. It contains 1,010 cases with clinical thoracic CT scans, of which 875 have lung nodule annotations acquired through a two-phase image annotation process performed by four experienced thoracic radiologists. To obtain a single ground truth, a 50% consensus criterion [24] is adopted to manage the variability among different radiologists. Since existing automatic segmentation algorithms cannot stably annotate lung regions in high-contrast CT scans, we excluded 509 cases with high-contrast CT scans and constructed the final dataset using the remaining 366 cases. All of these 366 cases serve as another validation cohort in our dataset organization, as shown in Table 2.
@@ 3.2.1 Evaluation metrics
In our experiments, we consider three image quality metrics: Fréchet Inception Distance (FID) [25], Maximum Mean Discrepancy (MMD) [26], and Mean Squared Error (MSE). FID is a commonly used metric to measure the distance between the distributions of real and synthetic data, as represented by deep learning features. We follow the implementation of HA-GAN [15] and utilize a 3D ResNet pretrained on a series of radiological datasets [27] to extract features from CT scans. MMD measures the difference between real and synthetic data distributions by comparing their moments in a high-dimensional feature space. We use the same 3D pretrained model to extract features for calculating MMD.
@@ 3.2.2 Implementation details
Table 4: Experimental results for the lung nodule segmentation evaluation                                                                                                                                                                                                                                                                         (The best evaluation score is marked in bold. $\uparrow$ means higher number is better, and $\downarrow$ indicates lower number is better. R. indicates real samples, and S. indicates synthetic samples.)
Discovery cohort (downstream) Method Dice ($\uparrow$) Sensitivity ($\uparrow$) Specificity ($\uparrow$)
800 R.: 100% discovery cohort (upstream) N/A 0.3598±0.0976 0.3704±0.1094 0.9999±0.0000
800 R. + 183 S.: 100% discovery cohort (upstream) + 50% validation cohort (LIDC-IDRI) MAV [14] GEM-3D [16] Med-DDPM [12] Lung-DDPM (OURS) 0.3190±0.1594 0.2991±0.2047 0.3537±0.1003 0.3908±0.0452 0.3290±0.1688 0.3487±0.1398 0.4046±0.2032 0.4393±0.1386 0.9999±0.0000 0.9980±0.0051 0.9998±0.0002 0.9999±0.0001
800 R. + 400 S.: 100% discovery cohort (upstream) + 50% validation cohort (IUCPQ) MAV [14] GEM-3D [16] Med-DDPM [12] Lung-DDPM (OURS) 0.3196±0.2232 0.3852±0.0553 0.3804±0.0534 0.3866±0.0293 0.3380±0.2389 0.4030±0.0509 0.3835±0.0695 0.3917±0.0449 0.9999±0.0001 0.9999±0.0000 0.9999±0.0000 0.9999±0.0000
800 R. + 800 S.: 100% discovery cohort (upstream) + 100% validation cohort (IUCPQ) MAV [14] GEM-3D [16] Med-DDPM [12] Lung-DDPM (OURS) 0.3577±0.0255 0.3584±0.0682 0.3638±0.0530 0.3914±0.0282 0.3674±0.0430 0.3891±0.0638 0.3729±0.0589 0.4078±0.0519 0.9999±0.0000 0.9998±0.0003 0.9999±0.0000 0.9999±0.0000
@@ 4.1.3 Ablation studies
We observe that performance in both image quality and lung nodule segmentation evaluation decreases when we remove the AAS process, highlighting its considerable contribution (FID: $0.0047\rightarrow 0.0600$, MMD: $0.0070\rightarrow 0.6617$, Dice: $0.3914\rightarrow 0.3869$). When using only lung nodule semantic layouts as the condition, we observe an increased FID score ($0.0047\rightarrow 0.0013$). This is because the AAS process additionally blends real lung images, allowing the synthetic images to be closer to the real ones. However, using only lung nodule semantic layouts negatively impacts the Dice score in the downstream task (Dice: $0.3914\rightarrow 0.3827$). If we remove both the AAS process and use only lung nodule semantic layouts, the model performance decrease dramatically (FID: $0.0047\rightarrow 0.0399$, MMD: $0.0070\rightarrow 0.4242$, Dice: $0.3914\rightarrow 0.3471$).
@@ 4.1.5 Performance in data scarcity scenarios
The motivation for proposing Lung-DDPM is to improve lung nodule segmentation performance in data scarcity scenarios. To evaluate this, we reduce the scale of the discovery (upstream) cohort to 25%, 50%, and 75%. We then use these reduced discovery cohorts to train three different pretrained models of our proposed method. Next, we generate three synthetic datasets according to the IUCPQ validation cohort and use them as extra synthetic samples to train and evaluate three different lung nodule segmentation models. The performance of these models is reported in Figure 5b. We find that the DICE score is highest when using 25% of the discovery cohort to train the proposed method, while the best sensitivity is achieved using 75% of the discovery cohort. These experimental results highlight that the proposed method can perform even better when faced with data scarcity, demonstrating its potential for deployment in practical scenarios.
@@ 5 Discussion
How to effectively utilize synthetic samples is a critical topic worth discussing. Existing studies have concluded that an appropriate amount of synthetic samples can be beneficial to downstream task performance [36, 37]. However, when narrowing our focus to lung nodule segmentation, the answer becomes ambiguous due to a lack of evidence. In Figure 5a, we confirm that a mixed training set for the lung nodule segmentation model, consisting of 75% real samples and 25% synthetic samples, is the most reliable and stable configuration.
@@
We summarize the strengths of our study as follows:
  • We proposed a novel DDPM-based thoracic CT image synthesis method that effectively and practically addresses three existing issues faced by current generative models.
  • A large number of patients from multiple institutions were evaluated, with images carefully annotated by experienced medical staff.
  • A comprehensive experimental design includes image quality assessments, a downstream lung nodule segmentation evaluation, and ablation studies for various impact factors.
The limitations of this study are as follows:
  • We focused on thoracic CT images without contrast and could not utilize a significant number of high-contrast CT scans from the LIDC-IDRI dataset due to the absence of a stable method for extracting lung annotations from them.
  • We only evaluated the proposed method at a single resolution. However, higher image sizes could enhance synthetic image quality.

### v2

@@ 2.1 Workflow
In this section, we detail the proposed diffusion model-based thoracic image synthesis method. The overall workflow is illustrated in Figure 2 and comprises four stages:
@@
    Data collection stage: The workflow begins with the collection of thoracic CT scans from clinical practice to establish a case management database. At this stage, no annotations are applied to the cases.
@@
    Training stage: The collected data is annotated by experienced radiologists (for lung nodules) and automatic algorithms (for lung areas) to balance accuracy and efficiency. The resulting annotated dataset is then used to train the proposed Lung-DDPM model.
@@
    Downstream task stage: In this stage, the synthetic samples are combined with real data to create a real/synthetic mixed dataset, which is used to train a lung nodule segmentation model. Training with this mixed dataset is expected to improve performance compared to training with real data alone.
@@ 2.2 Training and sampling processes
where $m$ is the semantic layout, $z\sim\mathcal{N}(0,\mathbf{I})$, $\sigma_{t}=\sqrt{\beta_{t}}$, and $\oplus$ is the concatenating operation between the semantic layout $m$ and the intermediate latent $x_{t}$. However, the traditional sampling process has a significant drawback: its overdependence on detailed semantic layouts. As demonstrated in Figure 1, without guidance from semantic layouts for extra-pulmonary areas, this process may produce abnormal results with incorrect anatomical structures. In clinical practice, manually annotating every anatomical structure in a thoracic CT scan is impractical. Moreover, semi-automatic and automatic segmentation tools [18, 19] are not sufficiently reliable to provide consistent semantic layouts due to the diverse types and parameter settings of CT scanners [20].
@@
where $m_{e}=1-m_{l}$ serves to spatially separate lung regions from extra-pulmonary regions during the sampling process. Note that even though $m_{e}$ itself does not contain anatomical semantic details, it acts as a spatial mask that enables dynamic fusion between synthesized lung regions (guided by $m_{l}$) and extra-pulmonary regions sampled from a real reference CT image. Through the proposed AAS process, we can generate high-quality thoracic CT images guided by radiologist-provided semantic layouts while maintaining anatomically reasonable extra-pulmonary areas, which significantly impact overall image quality. The training and AAS pseudo-codes are depicted in Algorithms 1 and 2, respectively.
@@ 2.3 Network designs
3D U-Net: We build upon the well-established U-Net backbone used in the original DDPM framework [17], which integrates residual blocks (ResBlks) and attention blocks (AttBlks) to enhance 3D representation learning. In our work, we extend this design to volumetric medical imaging by adapting it for 3D thoracic CT synthesis. While the use of ResBlks and AttBlks aligns with prior architectures, their integration within a 3D denoising diffusion framework introduces unique challenges in spatial modeling and memory efficiency, which we address through careful architectural calibration. Specifically, we modified it to accept a noise matrix and the corresponding lung semantic layout, both with dimensions of $128^{3}$. The network input consists of a noised CT image and its corresponding semantic layouts. After passing through a convolution layer, the encoder path (the left part of the 3D U-Net) progressively reduces spatial dimensions while increasing the number of feature channels, enabling the network to learn hierarchical representations of thoracic CT images. The bottleneck (the middle part of the 3D U-Net) processes low-dimensional features to capture global context and relationships within the CT image representations. Finally, the decoder (the right part of the 3D U-Net) reconstructs the spatial resolution of the features, gradually transforming the compact representation back into a full-sized output. Simultaneously, it combines both low-level and high-level features from the encoder via skip connections to generate detailed CT images. The decoder output is the predicted noise $\epsilon^{\prime}$ for the input image.
@@ 2.4 Learning object
According to Med-DDPM [12], although both L1 and L2 losses are commonly used in DDPM literature, L2 loss (unnormalized MSE loss) is more sensitive to outliers and tends to produce noisier results. In our model design, we follow the findings of Med-DDPM and adopt L1 loss as the learning objective to measure the differences between the added noise $\epsilon$ and the predicted noise $\epsilon^{\prime}$:
@@ 3.1 Description of cohorts
We utilize an internal (Quebec Heart and Lung Institute Research Center (IUCPQ)) dataset and an external dataset (LIDC-IDRI [22]) in this study. The population characteristics of the IUCPQ dataset are summarized in the online supplementary material. Additionally, we provide the patient list from the LIDC-IDRI dataset used in this study on the project’s GitHub page.
Table 1: Organization of our datasets (Synth. indicates synthetic samples.)
Cohort size Usage stage Data type
Discovery (upstream) 800 Model training (Lung-DDPM) Real
Validation (IUCPQ) 800 Synthetic sample generation Real $\rightarrow$ Synthetic
Validation (LIDC-IDRI) 366 External synthetic sample generation Real $\rightarrow$ Synthetic
@@
Test 249 Segmentation performance evaluation Real
IUCPQ [^1] The project was approved by the Comité d’éthique de la recherche of the Institut universitaire de cardiologie et de pneumologie de Québec-Université Laval (IUCPQ-UL), protocol number 2022-3752, 22156, on November 16, 2021, under the direction of Dr. Philippe Joubert.: The internal dataset includes 1,849 patients who underwent lung cancer resection and were collected by the IUCPQ. Each case has an original CT image and the corresponding lung/lung nodule semantic layouts. Every case in the dataset is nodule-positive. Lung nodules were annotated by Dr. Philippe Joubert and Josée Bafaro. To reduce the workload of the annotation team, we acquired lung semantic layouts by applying a U-Net-based automatic segmentation algorithm [19]. We split the data into four cohorts, each serving a specific role (Table 1).
@@
LIDC-IDRI: To further validate the developed Lung-DDPM under public and transparent experimental conditions, we included the external LIDC-IDRI [22] dataset, developed through collaboration between the Lung Image Database Consortium (LIDC) and the Image Database Resource Initiative (IDRI). This dataset consists of diagnostic and lung cancer screening thoracic CT scans with marked-up annotated lesions. It contains 1,010 cases with clinical thoracic CT scans, of which 875 have lung nodule annotations acquired through a two-phase image annotation process performed by four experienced thoracic radiologists. To obtain a single ground truth, a 50% consensus criterion [24] is adopted to manage the variability among different radiologists. Since existing automatic segmentation algorithms cannot stably annotate lung regions in contrast-enhanced CT scans, we excluded 509 cases with contrast-enhanced CT scans and constructed the final dataset using the remaining 366 cases. This cohort was used only for synthetic data generation and validation, not for training Lung-DDPM, as shown in Table 1.
@@ 3.2.1 Evaluation metrics
In our experiments, we consider three image quality metrics: Fréchet Inception Distance (FID) [25], Maximum Mean Discrepancy (MMD) [26], and Mean Squared Error (MSE). FID measures the distance between the distributions of real and synthetic data, represented by deep learning features. We follow the implementation of HA-GAN [15] and utilize a 3D ResNet pretrained on radiological datasets [27] to extract features from CT scans. MMD compares real and synthetic data distributions by evaluating moment differences in a high-dimensional feature space using the same pretrained model. MSE is computed as $\text{MSE}(x_{\text{real}},x_{\text{syn}},m_{l})=\frac{1}{\sum_{i}m_{l}^{(i)}}\sum_{i=1}^{N}m_{l}^{(i)}\left(x_{\text{real}}^{(i)}-x_{\text{syn}}^{(i)}\right)^{2}$, where $x_{\text{real}}$ and $x_{\text{syn}}$ are voxel intensities of the real and generated images, and $m_{l}$ is the binary lung mask indicating pulmonary regions. This metric is used to assess structural consistency between paired real and synthetic images conditioned on the same semantic layout. Note that for the only unconditional generative model HA-GAN, we do not report MSE since there is no one-to-one correspondence between generated and real images, making pixel-wise comparison ill-defined.
@@ 3.2.2 Implementation details
Table 3: Experimental results for the lung nodule segmentation evaluation                                                                                           (The best and the second evaluation score are marked in bold and underline, respectively. $\uparrow$ means higher number is better, and $\downarrow$ indicates lower number is better. R. indicates real samples, and S. indicates synthetic samples.)
Discovery cohort
Additional
data source
Dice ($\uparrow$) Sensitivity ($\uparrow$) Specificity ($\uparrow$)
800 R.:
100% discovery cohort (upstream)
N/A 0.3598±0.0976 0.3704±0.1094 0.9999±0.0000
800 R. + 183 R. or S.:
100% discovery cohort (upstream) +
50% validation cohort (LIDC-IDRI)
R.
MAV [14]
GEM-3D [16]
Med-DDPM [12]
Lung-DDPM (OURS)
0.3571±0.0969
0.3190±0.1594
0.2991±0.2047
0.3537±0.1003
0.3908±0.0452
0.3737±0.1535
0.3290±0.1688
0.3487±0.1398
0.4046±0.2032
0.4393±0.1386
0.9999±0.0001
0.9999±0.0000
0.9980±0.0051
0.9998±0.0002
0.9999±0.0001
800 R. + 400 R. or S.:
100% discovery cohort (upstream) +
50% validation cohort (IUCPQ)
R.
MAV [14]
GEM-3D [16]
Med-DDPM [12]
Lung-DDPM (OURS)
0.4044±0.0269
0.3196±0.2232
0.3852±0.0553
0.3804±0.0534
0.3866±0.0293
0.4053±0.0345
0.3380±0.2389
0.4030±0.0509
0.3835±0.0695
0.3917±0.0449
0.9999±0.0000
0.9999±0.0001
0.9999±0.0000
0.9999±0.0000
0.9999±0.0000
800 R. + 800 R. or S.:
100% discovery cohort (upstream) +
100% validation cohort (IUCPQ)
R.
MAV [14]
GEM-3D [16]
Med-DDPM [12]
Lung-DDPM (OURS)
0.4178±0.0212
0.3577±0.0255
0.3584±0.0682
0.3638±0.0530
0.3914±0.0282
0.4377±0.0301
0.3674±0.0430
0.3891±0.0638
0.3729±0.0589
0.4078±0.0519
0.9999±0.0000
0.9999±0.0000
0.9998±0.0003
0.9999±0.0000
0.9999±0.0000
@@ 4.1.2 Lung nodule segmentation evaluation
Moreover, Table 3 shows that adding real samples does not always lead to improved performance. For instance, incorporating 183 real samples from the external LIDC-IDRI dataset slightly decreases the Dice score from 0.3598 to 0.3571, indicating potential domain mismatch between the internal and external data. In contrast, adding 183 synthetic samples from Lung-DDPM yields a significantly higher Dice score of 0.3908 and improves sensitivity to 0.4393. Although adding 800 real samples from the internal IUCPQ validation cohort leads to the highest Dice score (0.4178), Lung-DDPM synthetic data still achieves comparable performance (Dice: 0.3914), offers close sensitivity (0.4078 vs. 0.4377), and preserves near-perfect specificity. These findings highlight that high-quality synthetic data from Lung-DDPM can outperform or rival real data augmentation in multiple scenarios, reinforcing its value for training segmentation models in data-scarce or privacy-constrained clinical environments.
@@ 4.1.3 Ablation studies
We observe that performance in both image quality and lung nodule segmentation evaluation decreases when we remove the AAS process, highlighting its considerable contribution (FID: $0.0047\rightarrow 0.0600$, MMD: $0.0070\rightarrow 0.6617$, Dice: $0.3914\rightarrow 0.3869$). When using only lung nodule semantic layouts as the condition, we observe an increased FID score ($0.0047\rightarrow 0.0013$). Without LS, the model relies solely on the lung nodule layout and the reference case during sampling. This reduces variation in the lung structure and leads to generation outcomes that more closely resemble the reference domain in global image statistics, which FID is sensitive to. Nonetheless, the absence of LS diminishes the model’s ability to control lung morphology explicitly, leading to poorer performance in downstream nodule segmentation, as evidenced by the reduced Dice Coefficient and sensitivity. However, using only lung nodule semantic layouts negatively impacts the Dice score in the downstream task (Dice: $0.3914\rightarrow 0.3827$). If we remove both the AAS process and use only lung nodule semantic layouts, the model performance decrease dramatically (FID: $0.0047\rightarrow 0.0399$, MMD: $0.0070\rightarrow 0.4242$, Dice: $0.3914\rightarrow 0.3471$).
@@ 4.1.5 Performance in data scarcity scenarios
The motivation for proposing Lung-DDPM is to improve lung nodule segmentation performance in data-scarce settings. To evaluate this, we reduced the size of the discovery (upstream) cohort to 25%, 50%, and 75% to train three different Lung-DDPM generative models. We then used these pretrained models to generate synthetic samples according to the IUCPQ validation cohort. These synthetic samples, combined with real data, were used to train and evaluate three different lung nodule segmentation models. The results are reported in Figure 5b.
Interestingly, segmentation models trained with synthetic data generated by a Lung-DDPM model trained on only 25% of the real discovery cohort often achieve comparable or even superior performance to those using synthetic data generated from a model trained on the full 100% discovery cohort. This counterintuitive result can be explained by considering the distribution of cases in the real dataset. The full 800-sample discovery cohort includes many small, low-contrast, ambiguous, or inconsistently labeled nodules that are inherently difficult to model and annotate accurately. Training the generative model on the full dataset exposes it to this variability, which can introduce artifacts or noise into the learned distribution and degrade the quality of generated samples. In contrast, the reduced 25% subset may overrepresent clearer, larger, and more consistently labeled nodules that provide stronger, less ambiguous learning signals. This allows the generative model to learn a cleaner, more canonical distribution of nodule appearances. Consequently, synthetic samples generated from this model exhibit higher anatomical fidelity and more consistent structure, leading to better segmentation performance when used for augmentation. This finding highlights that, in data-scarce medical imaging scenarios where annotation quality varies significantly, carefully selecting high-quality training data for generative models can be more valuable than simply maximizing dataset size, and that high-fidelity synthetic data can effectively supplement real data to train robust segmentation models.
@@ 5 Discussion
How to effectively utilize synthetic samples is a critical topic worth discussing. Existing studies have concluded that an appropriate amount of synthetic samples can be beneficial to downstream task performance [37]. However, when narrowing our focus to lung nodule segmentation, the answer becomes ambiguous due to a lack of evidence. In Figure 5a, we confirm that a mixed training set for the lung nodule segmentation model, consisting of 75% real samples and 25% synthetic samples, is the most reliable and stable configuration.